Problem III (20 points) Consider the following Eulers equation tx – 2tx + 2x = 3t”, t > 0. 1. (5 points) Make a substitutiPLEASE ANSWER ALL PARTS OF THE QUESTION
Answer #1

Solution boven Eulers ear is, +2 -24x! + 2x = 3+ to Rewriting, lt²0² 2to+2)x= 3t²t o li let t = 8² or z= loget and Did 0 -WY D ] (0,-1)00,-2) 3 (D,-2) (2-1) CD,-2) PI-32ě - PT. - 320 so, = logt P.I. - 3 logt & 2 Hence the general soin is given b